OpenStudy (anonymous):

Find the derivative of f(x) = 7/x at x = 1.

geerky42 (geerky42):

HINT: \(\dfrac{7}{x} = 7x^{-1}\) You can apply power rule.

OpenStudy (anonymous):

What's the power rule? I've only learned about the formula [f(x + h) - f(x)]/h

geerky42 (geerky42):

wow, you teacher is cruel, then. Give me time to figure that out.

geerky42 (geerky42):

OK, so while taking derivative, we have \(f'(x) = \displaystyle\lim_{h\to0}\dfrac{\dfrac{7}{x+h} - \dfrac{7}{x}}{h}\) In numerator, you can find GCD, which is \(x(x+h)\), right? So you have \(f'(x) = \displaystyle\lim_{h\to0}\dfrac{\dfrac{7}{x+h}\cdot\dfrac{x}{x} - \dfrac{7}{x}\cdot\dfrac{x+h}{x+h}}{h} =\lim_{h\to0} \dfrac{\dfrac{7x-7(x+h)}{x(x+h)}}{h}\) \[=\lim_{h\to0}\dfrac{\dfrac{7x-7x-7h}{x(x+h)}}{h}=\lim_{h\to0}\dfrac{\dfrac{-7h}{x(x+h)}}{h}=\lim_{h\to0}\dfrac{-7h}{xh(x+h)} \\~\\ =\lim_{h\to0}\dfrac{-7\cancel h}{x\cancel h(x+h)} = \lim_{h\to0}\dfrac{-7}{x(x+h)} = \lim_{h\to0}\dfrac{-7}{x^2+xh}\] Now you can do direct substitution; \(\cdots = \dfrac{-7}{x^2+x(0)} = \dfrac{-7}{x^2}\) Now plug in 1 for x.

geerky42 (geerky42):

You will learn about power rule soon... it will save you a lot of works...

geerky42 (geerky42):

@EuroSpeeder

OpenStudy (anonymous):

Oh wow that makes so much sense! So: -7/(1)^2 -7/1 = -7?

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Right

OpenStudy (anonymous):

Thank you so much!

geerky42 (geerky42):

no problem